OpenStudy (anonymous):

Rationalize the denominator of 5/(6 + 3sqrt2) - (2 + 2sqrt2)

OpenStudy (bahrom7893):

I'm assuming everything after / is in the denominator, so first let's simplify it a little.

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

To simplify it just expand the parenthesis and do some algebra to get: 6-2+3sqrt(2)-2sqrt(2) = 4 +sqrt(2)

OpenStudy (bahrom7893):

So now our fraction is just 5/(4+sqrt(2)) Do you know what rationalization does to a fraction?

OpenStudy (anonymous):

..no :/

OpenStudy (bahrom7893):

It gets rid of all the roots in the denominator. So how would you get rid of sqrt(2) in 4+sqrt(2)?

OpenStudy (bahrom7893):

I should say how would you get rid of the sqrt in 4+sqrt(2)?

OpenStudy (bahrom7893):

Hint: c/(a+b) = c*(a-b)/( (a+b)(a-b) )

OpenStudy (bahrom7893):

You have: 5/(4+sqrt(2))